OpenStudy (anonymous):

In order to maintain a healthy weight, an average adult dog that weighs 18 kilograms must consume 13 grams of dog food per kilogram of its body weight each day. A specialty pet store sells dog food by the cup, where 5 cups of dog food weighs 525 grams. About how many cups of dog food should the dog consume each day in order to maintain a healthy weight?

OpenStudy (anonymous):

the dosage given: 13 grams per kilogram of its body weight per day therefore for dog weighing 18 kilograms, it should be\[\frac{13~~grams}{\bcancel{kilograms}}\times18\bcancel{kilograms}=234~~grams~~per~~day\]to maintain a healthy weight... how many grams of dog food a pet store must provide based on the given information 5 cups = 525 grams?\[\frac{525~~grams}{5~~cups}=105~~grams~~per~cup\]to find the number of cups per day, we have to divide the number of grams per day by the number of grams per cap...\[\frac{234~~\bcancel{grams}~~per~~day}{105~~\bcancel{grams}~~per~cup}\approx2.23~~cups~~per~~day\]